OpenStudy (tiffany_rhodes):

Given that the acceleration vector is: a(t)=(-16cos(4t))i + (-16sin(4t))j + (-1t)k the initial velocity is: v(0)=i+k, and the initial position vector is: r(0)=i+j+k, compute the velocity vector v(t) and position vector r(t)

OpenStudy (freckles):

a (acceleration) is the second derivative of r (position) a (acceleration) is the first derivative of v (velocity)

OpenStudy (freckles):

that is r''=a r'=v

OpenStudy (freckles):

so to find v you need to integrate a to find v and integrate that v to find r

OpenStudy (tiffany_rhodes):

okay so would v(t) = (-4sin(4t)+c1, 4cos(4t)+c2, (-t^2)/2 + c3) ?

OpenStudy (freckles):

ok and v(0)=(1,0,1)

OpenStudy (freckles):

so \[-4\sin(4 \cdot 0)+c_1=1 \\ 4\cos(4 \cdot 0)+c_2=0\\ \frac{-0^2}{2}+c_3=1\]

OpenStudy (tiffany_rhodes):

and now we just solve for c1,2, and 3?

OpenStudy (freckles):

yep

OpenStudy (tiffany_rhodes):

alright, thank you :)

OpenStudy (freckles):

so that will give you the velocity vector

OpenStudy (freckles):

you will need to integrate velocity to find position vector

OpenStudy (freckles):

then apply the condition for position r(0)=(1,1,1)

OpenStudy (freckles):

shout if you need more help on this